# Backdoor Attack on Vertical Federated Graph Neural Network Learning

### v1

@@ 4.1.2 Datasets
This paper uses three widely adopted public datasets to evaluate the performance of BVG, including Cora [37], Cora_ml [37], and Pubmed [38]. The basic dataset statistics are summarized in Table 1.
@@ 4.1.3 Performance Metrics
To evaluate the performance of the BVG method, we employ two common metrics: attack success rate (ASR) and main task accuracy (MTA) [39]. ASR measures the percentage of samples in the backdoor test set that are predicted as the target class by the poisoned model. On the other hand, MTA assesses the accuracy of the clean test set on the poisoned model.
Bottom Datasets MTA ASR
Model Baseline GF TECB BVG Baseline GF TECB BVG
GCN Cora $69.7\pm 2.1$ 48.6$\pm$2.9 62.7$\pm$4.7 65.3$\pm$1.5 8.4$\pm$3.7 51.4$\pm$5.6 16.4$\pm$10.9 $99.9\pm 0.3$
Cora_ml $79.5\pm 2.1$ 64.2$\pm$7.6 77.1$\pm$0.9 76.5$\pm$2.6 4.8$\pm$2.1 26.8$\pm$5.2 34.6$\pm$7.7 $100.0\pm 0.0$
Pubmed $71.7\pm 2.1$ 44.0$\pm$10.2 64.9$\pm$4.2 71.3$\pm$3.5 6.5$\pm$2.2 24.0$\pm$8.0 81.4$\pm$13.8 $100.0\pm 0.0$
GAT Cora $67.8\pm 2.9$ 46.6$\pm$3.9 40.1$\pm$4.1 63.7$\pm$2.2 3.6$\pm$1.2 15.8$\pm$6.3 56.8$\pm$35.5 $99.1\pm 1.1$
Cora_ml $79.5\pm 1.1$ 61.0$\pm$3.3 57.8$\pm$4.9 75.6$\pm$2.8 3.6$\pm$1.6 23.4$\pm$2.7 45.1$\pm$31.0 $99.9\pm 0.2$
Pubmed $74.8\pm 1.0$ 70.8$\pm$3.5 45.0$\pm$17.9 73.3$\pm$1.7 4.5$\pm$0.8 30.2$\pm$2.0 89.6$\pm$20.8 $100.0\pm 0.0$
SGC Cora $68.9\pm 2.3$ 49.0$\pm$3.0 36.7$\pm$30.0 65.3$\pm$2.1 6.6$\pm$5.9 51.0$\pm$4.2 12.2$\pm$13.3 $99.8\pm 0.4$
Cora_ml $79.5\pm 1.4$ 67.4$\pm$4.0 74.8$\pm$2.6 74.9$\pm$2.4 4.6$\pm$2.5 27.8$\pm$4.0 15.2$\pm$9.1 $100.0\pm 0.0$
Pubmed $72.1\pm 2.0$ 41.8$\pm$12.3 54.0$\pm$27.0 70.6$\pm$2.9 7.1$\pm$1.0 32.7$\pm$10.9 59.8$\pm$48.8 $99.9\pm 0.2$
Table 2: Performance comparison with other methods on three datasets and three bottom models (Standard Deviation Included, Best Results Highlighted in Bold.)
@@ 4.1.4 Comparison Benchmarks
Due to the limited research on backdoor attacks against VFGNN, we chose two attack methods from other scenarios and the baseline VFGNN (without backdoors) for comparison. To make the attack methods from other scenarios transferable to VFGNN, we modified the comparison methods to a certain extent to match the VFGNN scenario and tasks. The comparison methods are briefly described as follows:
  • GF [36]: This method is designed for adversarial tasks on VFGNN. The original adversarial goal is to maximize the loss of node classification. To match our attack goal, we modified the adversarial goal to minimize the loss of the backdoor class. This method is mainly divided into three parts: obtaining the embedding vectors of other participants through queries, training a proxy model using the obtained embeddings, and obtaining the adversarial graph on the proxy model.
  • TECB [11]: This method is designed for backdoor attacks in VFL, mainly targeting image data. To adapt it to backdoor tasks on graphs, we set the target node’s attributes as the goal of its trigger embedding. This method is mainly divided into two stages: first, using limited known labels to train the trigger, and then using the trigger for target gradient alignment.
@@ 4.2 Experiment Results Analysis
4.2.1 Attack Efficacy in Two-Party Setting
We first conducted experiments in a VFGNN framework with two parties, and the obtained experimental results are shown in Table 2. To evaluate the impact of the BVG method on the main classification task, we also tested the performance of VFGNN without any backdoor attack (the Baseline in Table 2). The table shows that both comparison methods and BVG have a decline in the main task performance compared to the baseline, but BVG has the smallest decline, thus achieving the highest MTA among the attack methods. Table 2 also shows that BVG has a significantly higher backdoor attack success rate than those of to the compared methods for all the tested GNN models and used datasets. These results indicate that BVG achieves a high attack success rate with the least impact on the main task performance, making it an effective backdoor attack method to VFGNN.
4.2.2 Attack Efficacy in Multi-Party Setting
We also evaluated the proposed attack method’s performance in multi-party VFGNN scenarios where multiple participants are involved. Unlike the two-party setting, multi-party scenarios become more complex due to the increased number of participants and their interactions. Our evaluation focuses on how the presence of multiple participants affects the attack success rate and the main task performance of VFGNN.
@@ 4.2.2 Attack Efficacy in Multi-Party Setting
Figure 4: Evaluation of BVG under Defenses: The ISO, and GC Methods Demonstrate Protection Performance Against BVG with Various Noise Scales (0.5, 0.1, 0.01, 0.001) and Compression Rates (0.75, 0.5, 0.25, 0.1). DP Noise Scales are (0.75, 0.5, 0.25, 0.1). Max Norm Methods are Depicted as Single Points. The Optimal Position for a Defense Method is the Lower Right Corner, Indicating the Highest MTA and the Lowest ASR.
@@ 4.2.3 Attack Efficacy under Defense
To evaluate the robustness of the BVG method, we explored potential backdoor defense strategies within the VFGNN framework. We considered two main types of defense: the first type is GNN backdoor defense methods, such as Prune and Prune+LD [24]. These methods defend against backdoor attacks by pruning the edges between nodes with low similarity. Pruning these edges can disrupt the attacker’s trigger structure and connections. However, VFGNN’s architecture prevents defenders from pruning the adversary’s data, making these methods unsuitable. The second type is VFL backdoor defense methods, which reduce the leakage of labels, features, and other information during training by perturbing the gradients during the VFGNN training process, thereby lowering the likelihood of attackers implanting backdoors.
We implemented several representative VFL backdoor defense methods within the VFGNN framework. These defenses include Differential Privacy (DP) [18], max norm [40], Gradient Compression (GC) [18, 41], and adding Gaussian noise to intermediate features. These methods defend against backdoor attacks by perturbing the gradients sent from the active party to the adversary during VFGNN training. In our experiments, we set the compression rates of GC and the scale of DP to 0.75, 0.5, 0.25, and 0.1, and the scale of Gaussian noise to 0.5, 0.1, 0.01, and 0.001.
The obtained results are depicted in Figure 4 and show a trade-off between model utility and backdoor protection in VFGNN. Although different defense strategies can partially reduce the impact of backdoor tasks, they also degrade the performance of the main task. In our experiments with three underlying models and three datasets, ISO provided a relatively effective defense at a noise scale of 1, reducing the backdoor effect across all three datasets but at the cost of significantly lowering the accuracy of the main task. Therefore, the results indicate that typical defense methods cannot easily mitigate BVG attacks without severely sacrificing the performance of the main task, validating the robustness of the BVG method in resisting VFGNN defenses.
4.3 Analysis of Hyperparameter in BVG
4.3.1 Impact of Multi-hop Trigger in BVG

### v2

@@ 4.1.2 Datasets
This paper uses three widely adopted public datasets to evaluate the performance of BVG, including Cora McCallum et al. (2000), Cora_ml McCallum et al. (2000), and Pubmed Sen et al. (2008). The basic dataset statistics are summarized in Table 1.
@@ 4.1.3 Performance Metrics
To evaluate the performance of the BVG method, we use three common metrics: Attack Success Rate (ASR), Main Task Accuracy (MTA) Li et al. (2022b), and Mean Squared Error (MSE). Among them, ASR and MTA are the most commonly used evaluation metrics for backdoor attack tasks. ASR measures the proportion of samples in the backdoor test set that are predicted as the target class by the poisoned model. On the other hand, MTA assesses the accuracy of the clean test set on the poisoned model. MSE is used to measure the difference in the output of the bottom GNN model before and after trigger injection. A smaller MSE indicates that the injected trigger is less detectable by the active party, implying better stealth.
Bottom Datasets MTA
Model GF TECB VILLAIN ET NFT NFR BVG
GCN Cora 48.60$\pm$2.87 62.72$\pm$4.67 65.38$\pm$2.99 65.28$\pm$2.37 $66.12\pm 2.13$ 50.34$\pm$16.06 65.26$\pm$1.50
Cora_ml 64.20$\pm$7.63 $77.06\pm 0.91$ 73.60$\pm$3.49 75.50$\pm$2.78 75.46$\pm$2.25 56.90$\pm$21.80 76.50$\pm$2.63
Pubmed 44.00$\pm$10.20 64.90$\pm$4.15 69.26$\pm$2.86 59.80$\pm$15.40 68.44$\pm$2.58 32.36$\pm$13.57 $71.34\pm 3.54$
GAT Cora 46.60$\pm$3.93 40.06$\pm$4.10 63.44$\pm$2.01 61.70$\pm$6.37 $64.30\pm 3.04$ 41.02$\pm$16.66 63.72$\pm$2.20
Cora_ml 61.00$\pm$3.29 57.78$\pm$4.86 73.14$\pm$1.75 76.12$\pm$2.45 $76.50\pm 2.12$ 59.94$\pm$23.23 75.58$\pm$2.77
Pubmed 70.80$\pm$3.54 44.98$\pm$17.92 72.40$\pm$0.89 63.66$\pm$11.79 72.88$\pm$1.49 35.64$\pm$14.71 $73.28\pm 1.73$
SGC Cora 49.00$\pm$3.03 36.68$\pm$29.97 62.56$\pm$2.12 55.04$\pm$16.96 62.08$\pm$2.63 50.12$\pm$15.24 $65.30\pm 2.08$
Cora_ml 67.40$\pm$4.03 74.76$\pm$2.62 74.08$\pm$2.47 $75.34\pm 1.62$ 75.08$\pm$3.59 56.80$\pm$21.77 74.94$\pm$2.39
Pubmed 41.82$\pm$12.33 54.04$\pm$27.02 68.90$\pm$4.07 61.46$\pm$14.75 69.80$\pm$1.94 39.02$\pm$11.83 $70.64\pm 2.95$
Bottom Datasets ASR
Model GF TECB VILLAIN ET NFT NFR BVG
GCN Cora 51.40$\pm$5.61 16.40$\pm$10.91 99.00$\pm$0.64 48.80$\pm$13.62 91.54$\pm$4.13 36.48$\pm$22.01 $99.86\pm 0.28$
Cora_ml 26.80$\pm$5.15 34.60$\pm$7.74 97.34$\pm$3.94 49.92$\pm$7.82 74.78$\pm$22.86 38.12$\pm$31.24 $99.98\pm 0.04$
Pubmed 24.00$\pm$8.00 81.40$\pm$13.83 94.96$\pm$4.44 78.96$\pm$8.91 88.02$\pm$6.41 81.86$\pm$20.90 $100.00\pm 0.00$
GAT Cora 15.80$\pm$6.31 56.80$\pm$35.54 94.52$\pm$5.64 52.00$\pm$14.36 66.14$\pm$28.01 43.34$\pm$30.03 $99.12\pm 1.10$
Cora_ml 23.40$\pm$2.73 45.14$\pm$31.03 90.48$\pm$6.33 44.80$\pm$17.86 95.28$\pm$4.82 34.62$\pm$32.96 $99.92\pm 0.16$
Pubmed 30.20$\pm$2.04 89.60$\pm$20.80 86.98$\pm$10.71 78.42$\pm$9.90 98.64$\pm$2.23 80.62$\pm$20.43 $100.00\pm 0.00$
SGC Cora 51.00$\pm$4.20 12.24$\pm$13.28 99.34$\pm$0.65 54.40$\pm$22.44 87.60$\pm$13.96 32.46$\pm$19.15 $99.76\pm 0.39$
Cora_ml 27.80$\pm$3.97 15.18$\pm$9.14 98.72$\pm$1.42 60.22$\pm$12.20 88.66$\pm$7.56 37.78$\pm$31.25 $99.98\pm 0.04$
Pubmed 32.73$\pm$10.91 59.78$\pm$48.81 99.10$\pm$0.94 82.56$\pm$5.47 91.30$\pm$6.20 71.84$\pm$15.77 $99.90\pm 0.20$
Table 2: Performance comparison with other methods on three datasets and three bottom models (Standard Deviation Included, Best Results Highlighted in Bold. GF is an adversarial attack method, where a lower MTA is desirable.)
@@ 4.1.4 Comparison Benchmarks
Due to the lack of research on backdoor attacks in VFGNN, we selected three attack methods from other scenarios for comparison, adapting them to fit VFGNN’s tasks and structure. Additionally, we incorporated common GNN backdoor trigger schemes for further comparison. The methods are summarized below:
GF Chen et al. (2022): Originally designed for adversarial tasks in VFGNN, this method was adapted to minimize the loss of the backdoor class. It involves querying participant embeddings, training a proxy model, and generating an adversarial graph on the proxy model.
TECB Chen et al. (2023a): A VFL backdoor attack method for image data, adapted here for graph tasks by aligning target node attributes with trigger embeddings. It operates in two stages: training triggers with limited labels and aligning target gradients using the trigger.
VILLAIN Bai et al. (2023): Designed for VFL backdoor attacks, this method injects triggers into the bottom model’s output $H$, supporting clean-label attacks. It combines label inference using limited labeled samples and data poisoning with trigger masking and randomness for enhanced robustness.
Edge Triggering (ET): This method connects attack nodes to target class nodes via edges, creating a trigger relationship that exploits the graph’s topology to influence target node classification.
Node Feature Triggering (NFT): Backdoor attacks are achieved by injecting trigger patterns into nodes with known labels, altering their features to misclassify attack nodes as the target class under specific triggers.
Node Feature Replacement (NFR): Attackers replace the features of attack nodes with those of target class nodes, making their features similar to the target class and inducing misclassification.
@@ 4.2 Experiment Results Analysis
4.2.1 Backdoor Task Evaluation
We conducted experiments in a VFGNN framework with two participants, as shown in Table 2. BVG and NFT perform best on the main task, with slight variations across datasets and models. For the backdoor task, BVG consistently achieves the highest attack success rate across all tested models and datasets, demonstrating its effectiveness with minimal impact on main task performance.
We also evaluated the proposed attack method’s performance in multi-party VFGNN scenarios where multiple participants are involved. Unlike the two-party setting, multi-party scenarios become more complex due to the increased number of participants and their interactions. Our evaluation focuses on how the presence of multiple participants affects the ASR and the MTA of BVG.
@@ 4.2.1 Backdoor Task Evaluation
4.2.2 Backdoor Stealthiness Evaluation
Figure 4: Comparison of Backdoor Task Performance and Intermediate Layer Feature Differences Across Methods
In evaluating the stealthiness of backdoors in VFGNN, we measure the difference in the intermediate layer features of the same node before and after the backdoor injection. If the intermediate layer features show little change after the backdoor injection, we consider the backdoor difficult for the attacker to detect. For comparison, we selected several methods from Table 2 that perform relatively well on the backdoor task, with a particular focus on the performance of the BVG method under multi-hop triggers. Figure 4 shows the backdoor task performance and the average difference in intermediate layer features for five methods across three datasets.
From Figure 4, it is evident that the increase in MSE correlates positively with the increase in ASR. However, despite BVG achieving a significantly higher ASR than VILLAIN, its MSE is much lower than VILLAIN, indicating that BVG exhibits better stealthiness. For the other methods, BVG has slightly worse stealthiness but significantly better ASR performance. More importantly, the ASR performance of ET, NFT, and NFR has almost plateaued, as discussed in Appendix A.2.
@@ 4.2.3 Attack Efficacy under Defense
To evaluate the robustness of the BVG method, we explored potential backdoor defense strategies within the VFGNN framework. We considered two main types of defense: the first type is GNN backdoor defense methods, such as Prune and Prune+LD Dai et al. (2023). These methods defend against backdoor attacks by pruning the edges between nodes with low similarity. Pruning these edges can disrupt the attacker’s trigger structure and connections. However, VFGNN’s architecture prevents defenders from pruning the adversary’s data, making these methods unsuitable.
The second type of defense methods is specific to the VFL framework and can be divided into two categories: label inference defense and backdoor attack defense. Label inference defense operates on the assumption that backdoor attacks often require access to a large amount of label information. To prevent label leakage, defense methods perturb or compress gradients to minimize information leakage. Two representative methods are DP-SGD and gradient compression (GC). DP-SGD provides differential privacy protection by adding noise to gradients Fu et al. (2022), where the larger the noise amplitude, the smaller the privacy leakage, and vice versa. The GC method Fu et al. (2022); Kairouz et al. (2021) sends only a subset of gradients with the largest absolute values to participants, thereby reducing privacy leakage. For backdoor attack defense, there are mainly two approaches: one is model reconstruction, which aims to purify the top model of the active party to remove the backdoor, with methods such as model pruning Liu et al. (2018) and adversarial neuron pruning (ANP) Wu and Wang (2021). The other approach is to introduce interference into intermediate-layer features Li et al. (2021), such as embedding Gaussian noise into the intermediate features provided to the passive party to disrupt potential backdoor triggers.
Figure 5: The performance of BVG under DP-SGD defense
Figure 6: The performance of BVG under GC defense
In the VFGNN framework, we implemented several VFL backdoor defense methods, each with five parameter settings, as shown in the figure. We present the experimental results when the base model is GCN, with additional results provided in Appendix C. For label inference defense, as shown in Figures 5 and 6, our method is minimally affected, as the BVG method itself does not rely on label inference. For backdoor attack defense, the results of the ANP method are shown in Figure 7. The ANP method primarily affects model accuracy without effectively suppressing the backdoor. For the defense method involving interference with intermediate-layer features, the experimental results are shown in Figure 8. As model accuracy decreases, the success rate of backdoor attacks also drops, making it difficult for defenders to reduce ASR while maintaining a high MTA.
Figure 7: The performance of BVG under ANP defense
Figure 8: The performance of BVG under ISO defense
5 Conclusion
In this paper, we proposed the BVG method for ckdoor attacks to Vertical Federated Graph Neural Network (VFGNN). Utilizing a multi-hop trigger generation approach, the BVG method can perform efficient backdoor attacks with very limited knowledge of target class nodes. Extensive experiments on three datasets with three GNN models demonstrate that BVG achieves a high attack success rate while having a minimal impact on the main task accuracy and is unlikely to be detected by the active party. The evaluation of BVG efficacy under various defense methods highlights the robustness and efficiency of the proposed attack method, underscoring the necessity for advanced defense mechanisms in practical federated learning applications to counter such sophisticated backdoor attacks.
Appendix A Experimental Details
Our experiments are conducted in a VFGNN framework with only two participants by default. One is the active party, responsible for providing labels, and the other is the passive party, which provides data to assist the active party in training. Compared to Horizontal Federated Learning (HFL), the number of participants in Vertical Federated Learning (VFL) is typically smaller. This is because VFL requires the participants to have a large number of overlapping sample IDs while their features are complementary. However, in practice, it is almost impossible for dozens of companies with nearly identical user groups to simultaneously agree to collaborate. Following the experimental setup in existing VFL research, we assume there are only two participants by default Fu et al. [2022]; Bai et al. [2023].
For nine groups of experiments involving three datasets and three base models, we set the batch size for each training batch to 64, and the trigger update magnitude to 0.2. For the six experiments conducted on Cora_ml and Pubmed, we set the maximum trigger value $\epsilon$ to 0.08. For the experiments on Cora, we set the maximum $\epsilon$ to 0.1 for GCN and 0.08 for GAT and SGC.
In the Backdoor Retention process, we first train the model normally for the first five epochs, and then implement Backdoor Retention starting from the sixth epoch. To evaluate the backdoor effect, we randomly select 10 points for testing, noting that these 10 points do not need to be labeled points. In the experiments using GCN as the bottom model, the thresholds for the Cora, Cora_ml, and Pubmed datasets are set to 0.985, 0.98, and 0.98, respectively. In the experiments with GAT as the bottom model, the thresholds for the three datasets are set to 0.95, 0.98, and 0.98, respectively. For the SGC-based experiments, the thresholds for Cora, Cora_ml, and Pubmed are set to 0.9, 0.96, and 0.98, respectively.
This translation keeps the technical details intact while ensuring clarity in academic writing.
The experiments were run on a server equipped with an Intel(R) Xeon(R) Silver 4310 CPU @ 2.10GHz and an NVIDIA GTX 3090 GPU. Each experiment in Table 2 was conducted five times to evaluate the overall performance, and we recorded the average and standard deviation.
A.1 Design of Comparison Methods
To the best of our knowledge, there are currently no backdoor attack methods specifically designed for VFGNN. Therefore, in our research, we had to adopt some existing backdoor attack methods to evaluate the security of VFGNN. However, these methods are not entirely suitable for VFGNN, so we made appropriate adjustments to them. Below, we provide a detailed description of these adjustments.
GF Chen et al. [2022] is an adversarial attack method designed for VFGNN. In this method, the attacker first extracts global node embeddings from the system and constructs a shadow model that simulates the active party. Then, noise is added to the node embeddings, which are fed into the shadow model to explore vulnerabilities. Finally, adversarial samples are generated using the gradients derived from the shadow model. However, the objective of adversarial attacks is to misclassify the target node, whereas backdoor attacks aim to classify the target node into a specific class through certain trigger mechanisms. To adapt GF for backdoor attacks, we modified the final step by changing the optimization objective from minimizing the loss of an incorrect class to minimizing the loss of the backdoor class. All hyperparameters for GF remain consistent with those in the original paper.
TECB Chen et al. [2023a] is a backdoor attack method designed for VFL, with the trigger applied to the input samples. When the attacker intends to launch an attack, a trigger is added to the samples fed into the VFL model. However, this method was originally designed for image datasets, and its trigger pattern is not compatible with GNNs, which process graph-structured data. To adapt it to graph data, we modified the approach by setting the attributes of the target node to match the trigger embedding. TECB is configured with 4 known target samples, while all other hyperparameter settings remain consistent with the original paper.
VILLAIN Bai et al. [2023] is a backdoor attack method designed for VFL, where the trigger is applied to the output $H$ of the base model. This method assumes that the attacker must be a participant in the VFL (e.g., an institution) rather than a regular user, as ordinary users cannot modify the intermediate data exchanged between VFL participants. Despite the strong assumption, this design effectively avoids data type mismatches and can be directly applied to VFGNN scenarios. To evaluate its performance, we adjusted the hyperparameters of the original method, setting the trigger magnitude $\beta$ to 4, which is 10 times the default value in the original paper, to further enhance the attack’s effectiveness.
Edge Triggering (ET): Inspired by previous backdoor attack methods for GNNs Yang et al. [2022], we designed a method to inject backdoors by modifying the graph’s structure rather than node features. In this approach, the attacker connects the target node to known nodes of the target class via edges, forming a trigger relationship. This design leverages the GNN’s sensitivity to neighboring node features. Consistent with the BVG method, we assume the knowledge of 4 target class nodes. During training, these 4 nodes are interconnected, and during trigger injection, the target node is connected to these 4 nodes to carry out the backdoor attack.
Node Feature Triggering (NFT): Complementing the structure-based ET method, we designed a backdoor attack method that focuses solely on node features. Unlike the BVG method, where the trigger is optimized during training, we consider a fixed trigger pattern. A pre-designed trigger is injected into nodes with known labels, and the trigger pattern is consistent with that described in Bai et al. [2023]. To enhance the attack effect, we set the trigger magnitude $\beta$ to 4, we assume the knowledge of 4 target class nodes.
Node Feature Replacement (NFR): We further explore an extreme case where the attacker directly replaces the features of the target node with those of the target class node. This method reflects the sensitivity of VFGNN to node features. The feature replacement operation is expected to cause the target node to be misclassified into the target class, achieving a successful backdoor attack.
A.2 Performance of Comparison Methods
For the GF method, as it is fundamentally designed for adversarial tasks, its main task accuracy (MTA) essentially reflects its adversarial attack performance. In adversarial tasks, a lower MTA is desirable, which significantly differs from the goal of backdoor tasks—where maintaining a high MTA without affecting normal task performance is preferred. For this reason, we did not directly compare the MTA of the GF method in the paper, focusing instead on its attack success rate (ASR). From the perspective of ASR, the GF method demonstrates that adversarial attacks can effectively misclassify target nodes. In certain models and datasets, its ASR exceeds 50%. However, this result still falls short of achieving a practical backdoor attack, primarily because the adversarial performance of the GF method itself is suboptimal.
For the TECB method, we observed that its MTA and ASR are almost mutually exclusive: when the MTA is high, the ASR tends to be low, and vice versa. This characteristic makes TECB highly impractical in VFGNN scenarios. We believe the main reason for this is TECB’s second-stage poisoning of the top model. The original method targets image datasets, where the large parameter space and relative robustness of image data make poisoning effective for enhancing backdoor performance. However, in VFGNN, the smaller data and model parameter sizes significantly reduce the stability of the entire model during poisoning, leading to poor performance.
For the VILLAIN method, its backdoor task ASR and main task MTA both perform well, but overall, it is slightly inferior to the BVG method. Additionally, due to the design characteristics of its trigger mechanism, modifications to the intermediate feature $H$ are highly significant, greatly reducing the stealthiness of the backdoor. As shown in Figure 4, such noticeable feature changes make it challenging to evade detection in certain real-world scenarios.
For the ET method, as only minor adjustments to the graph topology were made during training, its impact on main task performance is minimal. However, the performance of its backdoor task is generally poor, with relatively good attack success only on the Pubmed dataset.
For the NFT method, its MTA performance is similar to that of the ET method. Since the trigger is injected into the feature of only one node at a time, its main task performance degradation is minimal. In terms of backdoor tasks, NFT demonstrates good effectiveness. However, it is worth noting that we set the trigger magnitude $\beta$ to 4, which is a very large value. We compared the trigger value range of NFT with that of BVG triggers and the input feature range (results are shown in Table 3). It can be seen that when $\beta$ is set to 4 for NFT, its trigger value range almost exceeds the normal range of input features, which could be considered illegal input in real-world scenarios.
For the NFR method, the arbitrary replacement of node features leads to a significant drop in main task performance, while its backdoor task performance is also poor. This indicates that the method struggles to maintain model stability while achieving effective attacks.
Bottom Datasets Origin BVG NFT
Model Data Trigger Trigger
GCN Cora 1.0000 0.1000 0.5916
Cora_ml 0.3623 0.0100 0.0982
Pubmed 0.0729 0.0100 0.2017
GAT Cora 1.0000 0.0800 0.5916
Cora_ml 0.3623 0.0100 0.0982
Pubmed 0.0729 0.0100 0.2017
SGC Cora 1.0000 0.0800 0.5916
Cora_ml 0.3623 0.0100 0.0982
Pubmed 0.0729 0.0100 0.2017
Table 3: Comparison of the Original Data Value Range with BVG Trigger and NFT Trigger Value Ranges
Appendix B Analysis of Hyperparameter in BVG
B.1 Impact of Multi-hop Trigger in BVG